OpenStudy (anonymous):

3(x + 1) = -2(x - 1) + 6

OpenStudy (kuoministers):

first thin you would do is to expand the brackets

OpenStudy (kuoministers):

thing*

OpenStudy (kuoministers):

after that you would collect like terms Then you can make x the subject to find x

OpenStudy (anonymous):

ooh okay then so what would it be thoe:)

OpenStudy (kuoministers):

ill do the first step for you then

OpenStudy (kuoministers):

3x + 3 = -2x +2 + 6

OpenStudy (anonymous):

okay then because i dont get it

OpenStudy (kuoministers):

do you get it now?

OpenStudy (kuoministers):

or shall i continue?

OpenStudy (anonymous):

no i need help on doing the problem

OpenStudy (anonymous):

i think u should continue

OpenStudy (kuoministers):

ok we need to collect like terms now. 5x = 5 now you can finish it off :)

OpenStudy (kuoministers):

everytime we take away something from one equation we need to do it to the other

OpenStudy (anonymous):

ooh okay then thank you so the answer is 25

OpenStudy (anonymous):

:)

OpenStudy (kuoministers):

actually no

OpenStudy (anonymous):

hold on so i got it wrong

OpenStudy (kuoministers):

5x = 5 we need to remove 5x to make it x so we divide 5x to make x so we need to divide the other side by 5 too \[\frac{ 5x }{ 5 } = \frac{ 5 }{ 5 }\]

OpenStudy (kuoministers):

so answer would be x = 1

OpenStudy (kuoministers):

do you get it? just read over what i typed. :)

OpenStudy (anonymous):

ooh i get it now i was doing it all wrong but thanks:)

OpenStudy (kuoministers):

If you think my answer helped you pls click best answer Thanks

OpenStudy (kuoministers):

Hmmm?